### v1

# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

Abstract
When we consider canonical averages for classical discrete systems, typically referred to as substitutional alloys, the map $\phi$ from many-body interatomic interactions to thermodynamic equilibrium configurations generally exhibits complicated nonlinearity. This canonical nonlinearity is fundamentally rooted in deviations of the discrete configurational density of states (CDOS) from continuous Gaussian families, and has conventionally been characterized by the Kullback-Leibler (KL) divergence on discrete statistical manifold. Thus, the previous works inevitablly missed intrinsic nonlinearities induced by discretization of Gaussian families, which remains invisible within conventional information-geometric descriptions. In the present work, we identify and quantify such unavoidable canonical nonlinearity by employing the 2-Wasserstein distance with a cost function aligned with the Fisher metric for Gaussian families. We derive an explicit expression for the Wasserstein distance in the limit of vanishing discretization scale $d\to 0$: $W_{2}=d\sqrt{\textrm{Tr}\left(\Gamma^{-1}\right)/12}$, where $\Gamma$ denotes covariance matrix of the CDOS. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, with future potential applications to configurational thermodynamics.
I Introduction
Canonical averages play a central role in statistical thermodynamics, providing a fundamental link between microscopic interactions and thermodynamic equilibrium configurations. For classical discrete systems with $f$ structural degrees of freedom (SDFs) on a given lattice, such as substitutional alloys, this correspondence is expressed as
$\Braket{q_{p}}_{Z}=Z^{-1}\sum_{i}q_{p}^{(i)}\exp\left(-\beta U^{(i)}\right),$ (1)
where $\{q_{1},\cdots,q_{f}\}$ denotes a complete set of structural coordinates, $\Braket{\cdot}_{Z}$ the canonical average, $\beta$ the inverse temperature, and $Z=\sum_{i}\exp\left(-\beta U^{(i)}\right)$ the partition function, with the summation taken over all microscopic configurations $i$.
When complete orthonormal basis functions (such as the generalized Ising modelce ) are adopted for the coordinates, the potential energy of configuration $k$ is given by
$U^{(k)}=\sum_{j=1}^{f}C_{j}q_{j}^{(k)},$ (2)
where the expansion coefficients are given by inner products $C_{j}=\Braket{U|q_{j}}$, i.e., trace over possible configurations.
Introducing the vectors $\mathbf{Q}_{Z}=(\Braket{q_{1}}_{Z},\cdots,\Braket{q_{f}}_{Z})$ and $\mathbf{U}=(C_{1},\cdots,C_{f})$, the canonical average in Eq. (1) defines a map
$\phi:\mathbf{U}\mapsto\mathbf{Q}_{Z}.$ (3)
It is well known that this map is generally nonlinear. Only in exceptional cases where the configurational density of states (CDOS) forms a multivariate Gaussian distribution, $\phi$ reduces to a globally linear map ig . In realistic discrete systems, however, the CDOS inevitably deviates from Gaussian families due to the discrete nature of configuration space imposed by lattice constraints.
Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence: The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS. While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves. Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports.
In the present work, we identify this contribution as an unavoidable canonical nonlinearity (UCN). To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian.
The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone. The details are shown below.
II Concept and Derivation
Unavoidable Canonical Nonliearity
Here, we briefly clarify the concept of canonical nonlinearity and its unavoidable contribution, referred to as the unavoidable canonical nonlinearity (UCN). Let $P\left(q\right)$ denote the CDOS of a realistic discrete system, given as a discrete probability distribution. We also introduce a reference continuous Gaussian distribution $P_{c}\left(q\right)$, which shares the same mean vector and covariance matrix $\Gamma$ as the realistic CDOS. It is well known that when the CDOS is exactly given by the continuous Gaussian $P_{c}$, the canonical average map $\phi$ in Eq. (3) becomesem2
$^{\forall}{C_{j}},\ \mathbf{Q}_{Z}=\left(-\beta\Gamma\right)^{-1}\cdot\mathbf{U},$ (4)
which implies that $\phi$ reduces to a globally linear map. From this viewpoint, canonical nonlinearity originates from deviations of the realistic CDOS from Gaussian families.
Based on this observation, previous studies have quantified canonical nonlinearity by measuring the difference between $P$ and $P_{c}$ using the KL divergence. Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$. Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
While this framework successfully captures non-Gaussian features inherent to the realistic CDOS, it inevitably includes an additional contribution originating from the discretization of the continuous Gaussian distribution itself. This contribution cannot be attributed to intrinsic non-Gaussian feature of the CDOS. Instead, it reflects geometric distortion of probability mass induced by the discretization process.
We identify this contribution as the unavoidable canonical nonlinearity, UCN. By construction, the UCN arises solely from discretizing continuous Gaussian families onto a discrete configurational support. Consequently, this contribution is fundamentally inaccessible to conventional information-geometric approaches formulated on discrete statistical manifolds, which compare probability weights only on fixed discrete supports. To isolate and quantify the UCN, it is therefore necessary to compare a continuous Gaussian distribution with its discretized counterpart in a framework that satisfies the following requirements: (i) it captures not only probability weights but also the geometric rearrangement of probability mass induced by discretization, and (ii) it remains consistent with, or admits a reinterpretation within, the existing information-geometric description on statistical manifolds. These considerations naturally motivate the present use of optimal transport theory.
Wesserstein Distance Aligned with Fisher Metric
Optimal transport theory can provide a natural framework to compare probability distributions defined on different supports, e.g., continuous and discrete ones. Among various transport distances, we employ the 2-Wasserstein distance $W_{2}$, which is expected to be particularly suitable for the present purpose, since it can provide compatible description for Gaussian families on statistical manifold, through its appropriate definition of the cost function.
Let $Q$ and $R$ be probability measures on $\mathbb{R}^{f}$. Then the squared 2-Wasserstein distance is given byw2
$W_{2}^{2}\left(Q,R\right)=\inf_{\pi\in\Pi\left(Q,R\right)}\int c\left(x,y\right)\,d\pi\left(x,y\right),$ (5)
where $\Pi(P,Q)$ denotes the set of all couplings (or transport plan) with marginals $P$ and $Q$, and $c\left(x,y\right)$ denotes cost function, i.e., transport cost from $Q\left(x\right)$ to $R\left(y\right)$. To ensure the compatibility with statistical manifold, we introduce a cost function as the quadratic form weighted by the inverse covariance matrix $\Gamma^{-1}$ of Gaussian family, namely,
$c\left(x,y\right)=\left(x-y\right)^{\mathrm{T}}\Gamma^{-1}\left(x-y\right).$ (6)
This choice can be viewed as replacing the standard Euclidean squared distance $\|x-y\|^{2}$ in the definition of the conventional 2-Wasserstein distance by the squared distance induced by the Fisher metric on the Gaussian statistical manifold. When we consider Gaussian families, its landscape is completely determined by mean vector $\mu$ and covariance matrix $\Gamma$. Considering that (i) $\mu\in\mathbb{R}^{f}$ and $\Gamma\in\mathbf{M}_{f,f}\left(\mathbb{R}\right)$ and (ii) transport object is in $\mathbb{R}^{f}$, appropriate Fisher metric is associated with translations of the mean while fixing covariance matrix, which reduces to the constant metric tensor $\Gamma^{-1}$. In fact, under the definition of the cost function of Eq. (6), we see the following relationship for Gaussian families:
$^{\forall}\delta\mu\in\mathbb{R}^{f},\$$W^{2}_{2}\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)$ (7)
$=2D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right),$
under the restriction of fixed covariance matrix and only allowing translations of its mean.
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq,\ d\to 0.$ (8)
In the limit $d\to 0$, the optimal transport plan is expected to converge to a local projection that maps each point $q\in V_{k}$ to the representative point $q^{\prime}_{k}$, namely,
$W_{2}^{2}\left(P_{c},P_{d}\right)=\sum_{k}\int_{V_{k}}\left(q-q^{\prime}_{k}\right)^{\textrm{T}}\Gamma^{-1}\left(q-q^{\prime}_{k}\right)P_{c}\left(q\right)dq.$ (9)
When we introduce the following variable transform
$u=q-q^{\prime}_{k},\ ^{\forall}i:\ \left|u_{i}\right|\leq\frac{d}{2}$ (10)
with representative hypercubic as $V_{0}$, and taking Tailor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
$W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $\sum_{k}P_{c}\left(q_{k}\right)\int_{V_{0}}u^{\textrm{T}}\Gamma^{-1}udu$ (11)
$=$ $\left\{\sum_{k}P_{c}\left(q_{k}\right)\right\}\left\{\sum_{i}\left(\Gamma^{-1}\right)_{ii}\frac{d^{f+2}}{12}\right\}$
$=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
The last equation can be obtained since at $d\to 0$,
$\sum_{k}P_{c}\left(q_{k}\right)d^{f}=\int_{\mathbb{R}^{f}}P_{c}\left(q\right)dq=1.$ (12)
Remarkably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process. This result provides a quantitative measure of the UCN induced purely by discretization.
Information-Geometric Intepretation of $W_{2}$
We finally provide an information-geometric interpretation of the derived Wasserstein distance in Eq. (11), clarifying its equivalence to the KL divergence associated with expected parallel translations of Gaussian distributions. Under fixed $\Gamma$, the KL between the following two Gaussian distribution is exactly given bykl
$^{\forall}\delta\mu\in\mathbb{R}^{f},\ D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)=\frac{1}{2}\delta\mu\Gamma^{-1}\delta\mu.$
This relation certainly shows that under fixed $\Gamma$, the KL divergence provides a well-known quadratic form with the (block-diagonal part of) Fisher metric tensor $\Gamma^{-1}$ as discussed above. To derive the $W_{2}$ distance in Eq. (11), we perform discretization of Gaussian for $f$-dimensional hypercubic with side length $d$. At the side of statistical manifold, we first extend the translation magnitude of Gaussian $\delta\mu$ to random vector as an i.i.d. random vector with probability density $\rho$ given by the uniform distribution, namely,
$\rho\left(\delta\mu\right)=\textrm{Unif}\left[-\frac{d}{2},\frac{d}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (14)
Under this extended definition, we consider the following expectation:
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=\sum_{i,k}\left(\Gamma^{-1}\right)_{ik}\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right].$ (15)
Since $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]$ corresponds to $\left(i,k\right)$-component of the covariance matrix for $\rho\left(\delta\mu\right)$, from Eq. (14),
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]_{i\neq k}=0$
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{i}\right]=\int_{-d/2}^{d/2}\frac{1}{d}x^{2}dx=\frac{1}{12}d^{2},$ (16)
thereby
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$ (17)
We therefore obtain the important relationships:
$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
$=$ $2\cdot E_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
Therefore, the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
III Conclusions
In this work, we investigated the intrinsic canonical nonlinearity arises from discrepancy between continuous Gaussian families and their discretized counterparts, based on optimal transport and information geometry. By focusing on the vanishing discretization limit for 2-Wasserstein distance with its cost function aligned with Fisher metric for Gaussian families, we demonstrated that the residual $W_{2}$ distance is determined solely by the inverse covariance matrix and the discretization scale. Our central result exhibits that the derived $W_{2}$ distance admits a clear information-geometric interpretation: it is exactly equivalent to twice the Kullback-Leibler divergence averaged over random parallel translations of the Gaussian induced by the discretization. Conceptually, this can provide a precise intuition for what we term unavoidable canonical nonlinearity within the statistical manifold, which has been in principle invisible effects solely under the KL-divergence based evaluation on a discrete statistical manifold
IV Acknowledgement
[…]