### v2

# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

Abstract
When we consider canonical averages for classical discrete systems, typically referred to as substitutional alloys, the map $\phi$ from many-body interatomic interactions to thermodynamic equilibrium configurations generally exhibits complicated nonlinearity. This canonical nonlinearity is fundamentally rooted in deviations of the discrete configurational density of states (CDOS) from continuous Gaussian families, and has conventionally been characterized by the Kullback-Leibler (KL) divergence on discrete statistical manifold. Thus, the previous works inevitablly missed intrinsic nonlinearities induced by discretization of Gaussian families, which remains invisible within conventional information-geometric descriptions. In the present work, we identify and quantify such unavoidable canonical nonlinearity by employing the 2-Wasserstein distance with a cost function aligned with the Fisher metric for Gaussian families. We derive an explicit expression for the Wasserstein distance in the limit of vanishing discretization scale $d\to 0$: $W_{2}=d\sqrt{\textrm{Tr}\left(\Gamma^{-1}\right)/12}$, where $\Gamma$ denotes covariance matrix of the CDOS. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems. In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale induces infinitesimal parameter variations.
I Introduction
Canonical averages play a central role in statistical thermodynamics, providing a fundamental link between microscopic interactions and thermodynamic equilibrium configurations. For classical discrete systems with $f$ structural degrees of freedom (SDFs) on a given lattice, such as substitutional alloys, this correspondence is expressed as
$\Braket{q_{p}}_{Z}=Z^{-1}\sum_{i}q_{p}^{(i)}\exp\left(-\beta U^{(i)}\right),$ (1)
where $\{q_{1},\cdots,q_{f}\}$ denotes a complete set of structural coordinates, $\Braket{\cdot}_{Z}$ the canonical average, $\beta$ the inverse temperature, and $Z=\sum_{i}\exp\left(-\beta U^{(i)}\right)$ the partition function, with the summation taken over all microscopic configurations $i$.
When complete orthonormal basis functions (such as the generalized Ising modelce ) are adopted for the coordinates, the potential energy of configuration $k$ is given by
$U^{(k)}=\sum_{j=1}^{f}C_{j}q_{j}^{(k)},$ (2)
where the expansion coefficients are given by inner products $C_{j}=\Braket{U|q_{j}}$, i.e., trace over possible configurations.
Introducing the vectors $\mathbf{Q}_{Z}=(\Braket{q_{1}}_{Z},\cdots,\Braket{q_{f}}_{Z})$ and $\mathbf{U}=(C_{1},\cdots,C_{f})$, the canonical average in Eq. (1) defines a map
$\phi:\mathbf{U}\mapsto\mathbf{Q}_{Z}.$ (3)
It is well known that this map is generally nonlinear. Only in exceptional cases where the configurational density of states (CDOS) forms a multivariate Gaussian distribution, $\phi$ reduces to a globally linear map ig . In realistic discrete systems, however, the CDOS inevitably deviates from Gaussian families due to the discrete nature of configuration space imposed by lattice constraints.
Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence:klo The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS. While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves. Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports.
In the present work, we identify this contribution as an unavoidable canonical nonlinearity (UCN). To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
In the final part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest. The details are shown below.
II Concept and Derivation
Unavoidable Canonical Nonliearity
Here, we briefly clarify the concept of canonical nonlinearity and its unavoidable contribution, referred to as the unavoidable canonical nonlinearity (UCN). Let $P\left(q\right)$ denote the CDOS of a realistic discrete system, given as a discrete probability distribution. We also introduce a reference continuous Gaussian distribution $P_{c}\left(q\right)$, which shares the same mean vector and covariance matrix $\Gamma$ as the realistic CDOS. It is well known that when the CDOS is exactly given by the continuous Gaussian $P_{c}$, the canonical average map $\phi$ in Eq. (3) becomesem2
$^{\forall}{C_{j}},\ \mathbf{Q}_{Z}=\left(-\beta\Gamma\right)^{-1}\cdot\mathbf{U},$ (4)
which implies that $\phi$ reduces to a globally linear map. From this viewpoint, canonical nonlinearity originates from deviations of the realistic CDOS from Gaussian families.
Based on this observation, previous studies have quantified canonical nonlinearity by measuring the difference between $P$ and $P_{c}$ using the KL divergence. Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$. Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
While this framework successfully captures non-Gaussian features inherent to the realistic CDOS, it inevitably includes an additional contribution originating from the discretization of the continuous Gaussian distribution itself. This contribution cannot be attributed to intrinsic non-Gaussian feature of the CDOS. Instead, it reflects geometric distortion of probability mass induced by the discretization process.
We identify this contribution as the unavoidable canonical nonlinearity, UCN. By construction, the UCN arises solely from discretizing continuous Gaussian families onto a discrete configurational support. Consequently, this contribution is fundamentally inaccessible to conventional information-geometric approaches formulated on discrete statistical manifolds, which compare probability weights only on fixed discrete supports. To isolate and quantify the UCN, it is therefore necessary to compare a continuous Gaussian distribution with its discretized counterpart in a framework that satisfies the following requirements: (i) it captures not only probability weights but also the geometric rearrangement of probability mass induced by discretization, and (ii) it remains consistent with, or admits a reinterpretation within, the existing information-geometric description on statistical manifolds. These considerations naturally motivate the present use of optimal transport theory.
Wesserstein Distance Aligned with Fisher Metric
Optimal transport theory can provide a natural framework to compare probability distributions defined on different supports, e.g., continuous and discrete ones. Among various transport distances, we employ the 2-Wasserstein distance $W_{2}$, which is expected to be particularly suitable for the present purpose, since it can provide compatible description for Gaussian families on statistical manifold, through its appropriate definition of the cost function.
Let $Q$ and $R$ be probability measures on $\mathbb{R}^{f}$. Then the squared 2-Wasserstein distance is given byw2
$W_{2}^{2}\left(Q,R\right)=\inf_{\pi\in\Pi\left(Q,R\right)}\int c\left(x,y\right)\,d\pi\left(x,y\right),$ (5)
where $\Pi(P,Q)$ denotes the set of all couplings (or transport plan) with marginals $P$ and $Q$, and $c\left(x,y\right)$ denotes cost function, i.e., transport cost from $Q\left(x\right)$ to $R\left(y\right)$. To ensure the compatibility with statistical manifold, we introduce a cost function as the quadratic form weighted by the inverse covariance matrix $\Gamma^{-1}$ of Gaussian family, namely,
$c\left(x,y\right)=\left(x-y\right)^{\mathrm{T}}\Gamma^{-1}\left(x-y\right).$ (6)
This choice can be viewed as replacing the standard Euclidean squared distance $\|x-y\|^{2}$ in the definition of the conventional 2-Wasserstein distance by the squared distance induced by the Fisher metric on the Gaussian statistical manifold. Note here that the $W_{2}$ distance between two Gaussian of $g\left(\mu_{1},\Gamma\right)$ and $g\left(\mu_{2},\Gamma\right)$ under the Euclidean metric is known as $\left|\mu_{1}-\mu_{2}\right|^{2}$.we When we consider Gaussian families, its landscape is completely determined by mean vector $\mu$ and covariance matrix $\Gamma$. Considering that (i) $\mu\in\mathbb{R}^{f}$ and $\Gamma\in\mathbf{M}_{f,f}\left(\mathbb{R}\right)$ and (ii) transport object is in $\mathbb{R}^{f}$, appropriate Fisher metric is associated with translations of the mean while fixing covariance matrix, which reduces to the constant metric tensor $\Gamma^{-1}$.aig In fact, under the definition of the cost function of Eq. (6) with the restriction of fixed covariance matrix and only allowing translations of its mean, $W_{2}^{2}=2D$ holds for any translation, supporting the compatibility of the introduced cost function in Eq. (6).
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
In the limit $d\to 0$, the optimal transport plan is expected to converge to a local projection that maps each point $q\in V_{k}$ to the representative point $q^{\prime}_{k}$, namely,
$W_{2}^{2}\left(P_{c},P_{d}\right)=\sum_{k}\int_{V_{k}}\left(q-q^{\prime}_{k}\right)^{\textrm{T}}\Gamma^{-1}\left(q-q^{\prime}_{k}\right)P_{c}\left(q\right)dq.$ (8)
When we introduce the following variable transform
$u=q-q^{\prime}_{k},\ ^{\forall}i:\ \left|u_{i}\right|\leq\frac{d}{2}$ (9)
with representative hypercubic as $V_{0}$, and taking Tailor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
$W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $\sum_{k}P_{c}\left(q_{k}\right)\int_{V_{0}}u^{\textrm{T}}\Gamma^{-1}udu$ (10)
$=$ $\left\{\sum_{k}P_{c}\left(q_{k}\right)\right\}\left\{\sum_{i}\left(\Gamma^{-1}\right)_{ii}\frac{d^{f+2}}{12}\right\}$
$=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
The last equation can be obtained since at $d\to 0$,
$\sum_{k}P_{c}\left(q_{k}\right)d^{f}=\int_{\mathbb{R}^{f}}P_{c}\left(q\right)dq=1.$ (11)
Notably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process. This result provides a quantitative measure of the UCN induced purely by discretization.
Information-Geometric Intepretation of $W_{2}$
We then provide an information-geometric interpretation of the derived Wasserstein distance in Eq. (10), clarifying its equivalence to the KL divergence associated with expected parallel translations of Gaussian distributions. Under fixed $\Gamma$, the KL between the following two Gaussian distribution is exactly given bykl
$^{\forall}\delta\mu\in\mathbb{R}^{f},\ D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)=\frac{1}{2}\delta\mu\Gamma^{-1}\delta\mu.$
This relation certainly shows that under fixed $\Gamma$, the KL divergence provides a well-known quadratic form with the (block-diagonal part of) Fisher metric tensor $\Gamma^{-1}$ as discussed above. To derive the $W_{2}$ distance in Eq. (10), we perform discretization of Gaussian for $f$-dimensional hypercubic with side length $d$. At the side of statistical manifold, we first extend the translation magnitude of Gaussian $\delta\mu$ to an i.i.d. random vector with probability density $\rho$ given by the uniform distribution, namely,
$\rho\left(\delta\mu\right)=\textrm{Unif}\left[-\frac{d}{2},\frac{d}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (13)
Under this extended definition, we consider the following expectation:
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=\sum_{i,k}\left(\Gamma^{-1}\right)_{ik}\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right].$ (14)
Since $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]$ corresponds to $\left(i,k\right)$-component of the covariance matrix for $\rho\left(\delta\mu\right)$, from Eq. (13), we can write:
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]_{i\neq k}=0$
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{i}\right]=\int_{-d/2}^{d/2}\frac{1}{d}x^{2}dx=\frac{1}{12}d^{2},$ (15)
thereby
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$ (16)
We therefore obtain the important relationships:
$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right)$
$=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
Generalization of the $W_{2}$-KL Equivalence beyond Gaussian Families
In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continous Gaussian. Although these derivations was shown for Gaussian families, the underlying mathematical structure indicates its broader applicability. We here briefly address the conditions where such a correspondence can be generalized.
Let $S$ be a statistical submanifold for a certain set of continous probability functions, (wholely or partly) parameterized by coordinates of
$\xi=\left(\xi_{1},\cdots,\xi_{f}\right),$ (18)
where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem). Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families. We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant. More generally, our statement rely on the locally-constant character of $\Omega$ under inifitesimal variations in $\xi$. Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
$c\left(q,q^{\prime}\right)=\left(q-q^{\prime}\right)^{\textrm{T}}\Omega\left(q-q^{\prime}\right),$ (19)
which ensures consistency between the transport cost and the local information-geometric structure of $S$.
Then, we discretize the underlying continuous space for $P$, by using $f$-dimensional hypercubic with side length $d$, and consider its limit of $d\to 0$. At the same time, we assume that the uncertainty or variation of the parameters $\xi$, induced by the discretization scales $d$, can be given by
$\delta\xi=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
where $a>0$ denotes a scalar constant, typically characterizing the sensitivity of the parameters $\xi$ to the discretization scale. Then we interpreted $\delta\xi$ as the following i.i.d. random vector with its distribution taking $\rho$:
$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (21)
Note that the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
Under these preparations, let us recall that KL-divergence between two nearby distributions in $S$ admits the following expansion:
$D\left(P_{\xi+\delta\xi}:P_{\xi}\right)=\frac{1}{2}\delta\xi^{\textrm{T}}\Omega\delta\xi+o\left(\left\|\delta\xi\right\|^{2}\right),$ (22)
where in the case of Gaussian families with $\xi=\mu$, the r.h.s. is exactly given by up to the quadratic form. Now it is clear that when we take expectation of KL-divergence in Eq. (22) w.r.t. $\rho\left(\delta\xi\right)$, its leading order at $d\to 0$ takes
$\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right]=\frac{a^{2}d^{2}}{24}\textrm{Tr}\left(\Omega\right).$ (23)
For the side of optimal transport theory, $W_{2}$ distance can be obtained in the same way as Gaussian families, namely,
$\lim_{d\to 0}W_{2}^{2}=\frac{d^{2}}{12}\textrm{Tr}\left(\Omega\right).$ (24)
Thereby, we obtain a generalized $W_{2}$-KL equivalence within the leading order at $d\to 0$:
$W_{2}^{2}\left(P:P^{\prime}\right)\sim\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
The proportionality constant explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$, and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed. In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.
III Conclusions
In this work, we investigated the intrinsic canonical nonlinearity arises from discrepancy between continuous Gaussian families and their discretized counterparts, based on optimal transport and information geometry. By focusing on the vanishing discretization limit for 2-Wasserstein distance with its cost function aligned with Fisher metric for Gaussian families, we demonstrated that the residual $W_{2}$ distance is determined solely by the inverse covariance matrix and the discretization scale.
Our central result exhibits that the derived $W_{2}$ distance admits a clear information-geometric interpretation: it is exactly equivalent to twice the Kullback-Leibler divergence averaged over random parallel translations of the Gaussian induced by the discretization. Conceptually, this can provide a precise intuition for what we term unavoidable canonical nonlinearity within the statistical manifold, which has been in principle invisible effects solely under the KL-divergence based evaluation on a discrete statistical manifold.
Finally, the present results indicate that the leading-order correspondence between the discretization-induced 2-Wasserstein distance and the expectation of the KL divergence reflects a general transport-information-geometric structure associated with the statistical submanifolds of interest. The unavoidable canonical nonlinearity observed in continuous Gaussian families and their discretized counterparts should therefore be understood as a particularly transparent example of this more general correspondence, where the parameter variations admit a clear physical interpretation in terms of parallel translations of Gaussian distributions.
IV Acknowledgement
[…]